Find an equation for the plane that passes through the point (-7,10,8), is parallel to the line x= 6−t, y= 9 +t, z=−5+ 2t,and intersects the plane 2x+y−z=e^pi^sqrt2at a 30◦angle.



1
Expert's answer
2021-09-26T18:08:32-0400

The equation of the plane will be


"a(x-x_1)+b(y-y_1)+c(z-z_1)=0"

Given "x_1=-7, y_1=10, z_1=8."

"\vec n=\langle a, b, c \rangle"

The plane is parallel to the line "x= 6\u2212t, y= 9 +t, z=\u22125+ 2t"


"\vec n\cdot\langle -1, 1, 2\rangle=0"

"-a+b+2c=0"

The plane intersects the plane "2x+y−z=e\cdot\pi\cdot\sqrt{2}" at a "30\degree." angle.


"\cos 30\degree=\dfrac{2a+b-c}{\sqrt{a^2+b^2+c^2}\sqrt{4+1+1}}"

"9a^2+9b^2+9c^2=2(2a+b-c)^2"

Let "a=1." Then


"b=1-2c"

"9+9(1-2c)^2+9c^2=2(2+1-2c-c)^2"

"9(1+1-4c+4c^2+c^2)=2(9)(1-2c+c^2)"

"5c^2-4c+2=2c^2-4c+2"

"3c^2=0=>c=0"

"b=1"

Substitute


"1(x-(-7))+1(y-10)+0(z-8)=0""x+7+y-10=0"

The equation of the plane will be


"x+y=3"